OpenStudy (anonymous):

4+log subscript 9(3x-7)=6

OpenStudy (campbell_st):

subtract 4 from both sides \[\log _{9} (3x - 7) = 2\] raise each side as a power of 9 and remember \[9^{\log _{9}(3x - 7)} = 3x - 7\] then \[(3x - 7) = 9^2\] 3x - 7 = 81 3x = 88 I'll let you finish the problem

OpenStudy (anonymous):

did u get 261?

OpenStudy (phi):

where did you get 261 from?

OpenStudy (anonymous):

i subed the 88 in for x and did that whole problem over is that wrong??

OpenStudy (phi):

the problem is asking you to find an x, so that when you put it into the original equation, it will be true. campbell figured out the 3x= 88 that means 3 times x (the number we are looking for) is 88 or x+x+x= 88 x is definitely less than 88 (it takes 3 of them to get 88) you solve for x by dividing both sides by 3

OpenStudy (anonymous):

wow i feel slow -_-

OpenStudy (phi):

If this were easy it would not be math

OpenStudy (phi):

so with 3x=88 what is x?

OpenStudy (phi):

Do you know how to find x if 3x=88 ?

OpenStudy (anonymous):

no :/

OpenStudy (fellowroot):